# Deciding Subspace reachability problems with application to Skolem’s Problem

### v1

@@ 3. Proofs of Theorems
In Proposition 4, there is a single dominating characteristic root, while in the case of Theorem 3 there are $r\geq 1$ characteristic roots of maximal modulus. However, by assumption, there is a single real root, $\lambda_{1}$, whose algebraic multiplicity in the minimal polynomial $P_{A}$ of $A$ is larger than the algebraic multiplicity of the other dominating roots. This implies that, of the $r$ dominating roots, the Jordan block of largest dimension which is associated with $\lambda_{1}$ dictates the asymptotic behavior. Then by Lemma 6, after projecting to real projective space, orbits approach the projection of a line, and hence must eventually be bounded away from at least one of the $E_{i}$ after a finite number of iterations, where the $E_{i}$ are as defined in the proof of Proposition 4. ∎
@@ 4. Concluding remarks
This paper does not come close to reaching the limits of the techniques presented here, nor do we apply this machinery to every problem vulnerable to our methods. Indeed, the results of this paper indicate that deepening our understanding of dynamical systems in real projective space $\mathbb{R}\mathbb{P}^{d-1}$ with maps induced by matrices in $GL(d,\mathbb{R})$ provides insight into the higher-dimensional Orbit Problem, and more generally termination problems, by way of the arguments given in this paper. And, possibly, the geometric and dynamical mechanisms penetrating the Orbit Problem may link to the algebraic and number-theoretic structures traditionally employed in nontrivial ways. In particular, we believe that mixing the general geometric and dynamical structures provided here with the finer algebraic and number-theoretic tools traditionally used can lead to additional breakthroughs in this area.
@@
Next, we reflect that the essential reason why we require that $W\cap U=\{0\}$ in order to have decidability in Theorem 1, follows from the main observation this paper makes, which is that orbits approach the projection of $W$ in real projective space, i.e. the angles between orbits and $W$ monotonically approaches zero, and as a consequence whenever $W\cap U=\{0\}$ orbits can be bounded away from $U$ in finite time. Fortunately, if $A,x$, and $U$ are randomly generated by drawing entries at random from some finite set, when $A$ has $r$ dominating roots, $p$ of which have highest algebraic multiplicity in the minimal polynomial of $A$, and $U$ dimension $\leq d-p$, then with overwhelmingly high probability $W\cap U=\{0\}$. Thus, arguing from this intuitive level, Theorem 1 decides a “large” class of instances.

### v2

@@ 3. Proofs of Theorems
In Proposition 5, there is a single dominating characteristic root, while in the case of Theorem 3 there are $r\geq 1$ characteristic roots of maximal modulus. However, by assumption, there is a single real root, $\lambda_{1}$, whose multiplicity in the minimal polynomial $P_{A}$ of $A$ is larger than the multiplicity of the other dominating roots. This implies that, of the $r$ dominating roots, the Jordan block of largest dimension which is associated with $\lambda_{1}$ dictates the asymptotic behavior. Then by Lemma 6, after projecting to real projective space, orbits approach the projection of a line, and hence must eventually be bounded away from at least one of the $E_{i}$ after a finite number of iterations, where the $E_{i}$ are as defined in the proof of Proposition 5. ∎
We now move to consider the continuous Orbit Problem, studying flows rather than maps. It is in this setting that the foregoing development of dynamics in projective space is most useful: as in the discrete case, every flow is asymptotic to some stationary point or cycle. Then, we have a way of bounding the location of flows in projective space, which aids in decidability; particularly because when considering flows, traversal from one side of a boundary to another implies intersecting the boundary.
Proof of Theorem 4..
In the case $A$ has a complex-conjugate pair of eigenvalues, the projection of the flow onto projective space is simply a cycle. Hence the cycle necessarily intersects with the target line, which is a point in $\mathbb{RP}$, infinitely often. This holds independently of whether the real component of the complex eigenvalue pair is negative, positive, or zero.
Whenever $A$ has two distinct real eigenvalues or a single repeated real eigenvalue, the flow is stationary when the eigenvalue(s) are zero, and otherwise if not zero the flow approaches a stationary point $p\in\mathbb{RP}$, computable from $A$ by translating to the Jordan normal form. Since the target space $U$ has rational entries, it is decidable whether $p=\mathbb{P}U$. If $p=\mathbb{P}U$, and $x(0)\neq U$, then $x(t)\not\in U$ for all $t\in\mathbb{R}^{+}$, since $U$ is a stationary point in projective space and flows are unique. Otherwise, when $p\neq\mathbb{P}U$, check if $\mathbb{P}x(0)$ and $p$ are on the same side of $\mathbb{P}U$. If so, $x(t)\not\in U$ for all $t\in\mathbb{R}^{+}$. If not, there exists a $t\in\mathbb{R}^{+}$ such that $x(t)\in U$ because $\mathbb{P}x(t)$ is forced to cross $\mathbb{P}U$ on approach to $p$. ∎
@@ 4. Concluding remarks
This paper does reach the limits of the techniques presented here, nor do we apply this machinery to every problem vulnerable to our methods. Indeed, the results of this paper indicate that deepening our understanding of dynamical systems in real projective space $\mathbb{R}\mathbb{P}^{d-1}$ with maps induced by matrices in $GL(d,\mathbb{R})$ provides insight into the higher-dimensional Orbit Problem, and more generally termination problems, by way of the arguments given in this paper. And, possibly, the geometric and dynamical mechanisms penetrating the Orbit Problem may link to the algebraic and number-theoretic structures traditionally employed in nontrivial ways. In particular, we believe that mixing the general geometric and dynamical structures provided here with the finer algebraic and number-theoretic tools traditionally used can lead to additional breakthroughs in this area.
@@
Next, we reflect that the essential reason why we require that $W\cap U=\{0\}$ in order to have decidability in Theorem 1, follows from the main observation this paper makes, which is that orbits approach the projection of $W$ in real projective space, i.e. the angles between orbits and $W$ monotonically approaches zero, and as a consequence whenever $W\cap U=\{0\}$ orbits can be bounded away from $U$ in finite time. Fortunately, if $A,x$, and $U$ are randomly generated by drawing entries at random from some finite set, when $A$ has $r$ dominating roots, $p$ of which have highest multiplicity in the minimal polynomial of $A$, and $U$ dimension $\leq d-p$, then with overwhelmingly high probability $W\cap U=\{0\}$. Thus, arguing from this intuitive level, Theorem 1 decides a “large” class of instances.